A piston moves from rest to a speed of 5 m/s in one twenty-fifth of a second. What are the average acceleration and the distance travelled?


1
Expert's answer
2021-06-01T10:47:56-0400

"t =\frac{1}{25}=0.04"

"\vec v_0=0;\vec v_1=5"

"\vec a=\frac{\vec v_1-\vec v_0}{t}=\frac{5}{0.04}=125"

"\vec s=\vec v_0t+\frac{at^2}{2}=\frac{125*0.04^2}{2}=0.1"

"\text{Answer: }\vec a=125\frac{m}{s^2};\vec s=0.1m"